If 20%of x=y and 69%of(x+y) =360 then find x and y
With process please
 

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
20x /100 = y
x = 5y
 
Now
 
69(x+y)/100 = 360
 
Put x= 5y
6y = 360*100/69
y = 6000/69
 
Regards
Arun (askIITians forum expert)